OpenStudy (anonymous):

At a large nursery, a border for a rectangular garden is being built. Designers want the border's length to be 5 ft greater than its width. A maximum of 180 ft of fencing is available for the border. Write and solve an inequality that describes possible widths of the garden.

OpenStudy (anonymous):

\[2 (2 w+5)\leq 180 \]

OpenStudy (anonymous):

L=w+5; 2L+2w=180; 2(w+5)+2w=180; 4w+10=180 W=170/4 W=85/2... Same as robtobey